if the length and breadth of a rectangular field are increased, the area increases by 50%. if the increase in the length was 20% , by what percentage was the breadth increased?
Answer by swincher4391(1107)

.2l * xb = .5A
.2*x (l*b) = .5A
.2*x A = .5A
.2*x = .5
x = 2.5
It would have to be increased 250%.